OpenStudy (anonymous):

5x+2y=10 Solve for x then solve for y then graph

OpenStudy (igreen):

If you solve for x: 5x + 2y = 10 Subtract 2y to both sides: 5x = -2y + 10 Divide 5 to both sides of the equation: x = -2/5y + 2

OpenStudy (igreen):

If you solve for y: 5x + 2y = 10 Subtract 5x to both sides: 2y = -5x + 10 Divide 2 to both sides of the equation: y = -5/2x + 5

OpenStudy (anonymous):

Graph it

OpenStudy (igreen):

Graph

OpenStudy (igreen):

Does it want you to graph x = -2/5y + 2 and y = -5/2x + 5 too?

OpenStudy (igreen):

Well that's the graph for 5x + 2y = 10, x = -2/5y + 2, and y = -5/2x + 5. Because they're all the same, and you just changed the order of the values.

OpenStudy (morganchavez):

x=2-2yover5 y=5-5x over 2

OpenStudy (morganchavez):

http://www.mathway.com/problem/MzIxOTgyNDgy

OpenStudy (morganchavez):

thts for y

OpenStudy (morganchavez):

http://www.mathway.com/problem/MzIxOTgzMzUx

OpenStudy (morganchavez):

thts for x

OpenStudy (morganchavez):

http://www.mathway.com/problem/MzIxOTgzODk0

OpenStudy (morganchavez):

thts for both

OpenStudy (morganchavez):

@blove555